solve the linear system by using elimination
3x - 4y = -16
x - 4y = -40

Answer by ewatrrr(24785)

 
Hi
solve the linear system by using elimination
3x - 4y = -16
x - 4y = -40
3x - 4y = -16
-x + 4y = 40 |multiplying 2nd EQ by -1 and adding to 1st EQ
2x = 24
x = 12 and y = 13 |x-4y=+-40, 12+40 = 4y, y = 52/4
CHECKING our answer***
using: 3x - 4y = -16
36 - 52= -16
Wish You the Best in your Studies.